# Faster Cascades via Speculative Decoding

### v1

@@ 4.4 Optimal speculative deferral
One can now run the speculative cascading procedure in Algorithm 5 using (10) as the deferral rule; the corresponding $\delta(\cdot)$ is listed in Table 1. See §C.2 for a similar derivation for $\ell=\ell_{\log}$.
Remark 3 (Optimal deferral with temperature sampling).
In practice it is common to apply a temperature scaling to both $q_{t}$ and $p_{t}$, and sample from $\tilde{q}_{t}(v)\propto q_{t}(v)^{1/T}$ and $\tilde{p}_{t}(v)\propto p_{t}(v)^{1/T}$ respectively, for a temperature parameter $T>0$. In this case, the constrained problem in (7) would use the TV distance between the temperature-scaled distributions $\tilde{q}_{t}$ and $\tilde{p}_{t}$ to measure the resulting rejection rate, and the optimal deferral rule in Lemma 4 would now use $D_{\textup{{TV}}}(\tilde{p}_{t},\tilde{q}_{t})$ instead of $D_{\textup{{TV}}}({p}_{t},{q}_{t})$. To construct a plug-in estimator to this optimal rule, we still prescribe using the unscaled probabilities $q_{t}$ and $p_{t}$ to estimate the expected loss, giving us, for $\ell=\ell_{\emph{\text{0-1}}}$:
$\textstyle\tilde{r}_{\rm\tt OPT}(x_{<t})=1~{}~{}\iff~{}~{}\max_{v}q_{t}(v)\,<\,\max_{v}p_{t}(v)-\alpha\cdot D_{\textup{{TV}}}(\tilde{p}_{t},\tilde{q}_{t}).$
One may run Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule and the temperature scaled $\tilde{q}_{t}$ as the drafter.
Remark 4 (Special case of greedy decoding).
When temperature $T\rightarrow 0$, $D_{\textup{{TV}}}(\tilde{p}_{t},\tilde{q}_{t})=1$ whenever $\operatorname{argmax}_{v}p_{t}(v)\neq\operatorname{argmax}_{v}q_{t}(v)$, and is zero otherwise. In this case, running Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule (and $\tilde{q}_{t}$ as the drafter) is equivalent to running it with $\hat{r}_{\rm\tt Diff}$ in (5) as the deferral rule. In other words, for greedy decoding, the optimal deferral rules for a speculative cascade coincides with that for a sequential cascade. We formalize this in Lemma 8 in §C.3.
5 Further related work
@@
There has been a stream of work on improving the draft generation process in speculative decoding; these include having the drafter and verifier share the same backbone [39, 25, 6, 30, 21, 51, 16, 27], using multiple small draft models [9, 46], using tree-structured draft batches [38, 29], distilling the drafter with the verifier [52], and leveraging multiple sampled candidates from the drafter [40].
@@
The work that is most closely related to our specific proposal is the Big Little Decoder (BiLD) [24], which can be seen as another lossy variant of speculative decoding [26, 44, 52]. BiLD has two phases: a fallback phase, during which the drafter $q$ is run auto-regressively until its maximum predicted probability is sufficiently low; and a rollback phase, during which the verifier $p$ is run in parallel on the prefixes generated by $q$ and rolls back to the point where $D(q,p)>\alpha$, for a metric $D$ that measures discrepancy and threshold $\alpha$. The fallback phase can be seen as implementing Chow’s deferral rule (2), and allows for the draft window size to vary dynamically based on an estimate of how likely the draft tokens will be accepted; the rollback phase can be seen as a deterministic variant of the rejection sampling algorithm of [26].
@@
An advantage of BiLD over the rejection sampling algorithm in [26] is the use of Chow’s rule to vary the draft window size. However, the final target distribution it seeks to mimic, $\mathbb{T}_{\textup{BiLD}}(q,p)(v)=\mathbf{1}(D(q,p)\leq\alpha)\cdot q(v)+\mathbf{1}(D(q,p)>\alpha)\cdot p(v)$, is an approximation to $p$, in that the target distribution $\pi=\mathbb{T}_{\textup{BiLD}}(q,p)$ is chosen to satisfy $D(\pi,p)\leq\alpha$. Hence, in cases where $q$ deviates substantially from $p$, BiLD would choose $p$ as the target distribution, even $q$ offers better quality (as captured by a suitable loss function in §4.4) on a prefix. In contrast, our proposed approach in §4 uses speculative decoding to approximate target distributions that optimally cascade between $q$ and $p$. In our experiments, we compare the efficacy of using $\mathbb{T}_{\textup{BiLD}}$ as the target distribution with the target distributions we propose in this paper (see Table 1). propose in this paper (see Table 1).
6 Experimental results
We present experiments to demonstrate that the proposed speculative cascading techniques yield better cost-quality trade-offs compared to both sequential token-level cascades and standard speculative decoding. We construct speculative cascades with three different deferral rules: (i) Chow in (2), (ii) Diff in (5) and (iii) OPT in (10). Our experimental setup is based on [52]; see D.1 for details.
Tasks and models.
[…]
Baselines.
Since our primary goal is to demonstrate the benefits of combining both cascading and speculative decoding, we compare against representative methods from both paradigms:
  1. (i)
    Sequence-level cascade [22, 18] based on sequence-level Chow’s rule in (1) (SeqCascade [Chow])
  2. (ii)
    Token-level cascade outlined in Algorithm 2, with token-level Chow’s rule in (2) used for deferral [10, 19] (TokenCascade [Chow])
  3. (iii)
    Lossy speculative decoding described in §2, with both $\beta=1$ [26, 52] (SpecDecode [Lossy]) and with $\beta$ tuned using the procedure in [44] (SpecDecode [Lossy])
  4. (iv)
    A variant of the Big-Little Decoder approach [24], which applies Algorithm 4 to the target distribution $\mathbb{T}_{\textup{BiLD}}$ in §5 (BiLD).
We do not include the oracle approach in Algorithm 3 as it is impractical to evaluate its running time. We also note that, while one could potentially integrate other recent variants of the speculative decoding (e.g., those in §5 that warrant changes to the model architectures) into our speculative cascades framework, in the interest of a fair comparison, we use the original version by [26].
We evaluate all methods under both greedy decoding ($T=0$), and temperature sampling with temperature $T=1$. As noted in §4.3, with greedy decoding, the OPT deferral rule coincides with the Diff deferral rule. We set the block size $\gamma=5$ for the methods based on speculative execution.
Results.
In Figure 2, we present plots of quality versus latency for the different methods. In each case, we vary the lenience parameter $\alpha$, and plot either the BLEU or ROUGE-2 metric as a function of the relative latency to the larger model. For brevity, we include the three main baselines; in §D.3 we compare to SpecDecode [Lossy] [44].
Clearly, methods that use speculative execution are considerably faster than sequential token-level cascades, although sequential cascades do have a slight advantage in the low-latency regimes (unlike speculative approaches, which always call the large model after every $\gamma$ steps, a sequential cascade only invokes the large model when the small model defers). Under temperature sampling ($T=1$), the OPT speculative cascading strategy is seen to provide the best quality metrics for most latency values, with Diff coming in a close second. A similar trend is also seen with greedy decoding, where the optimal deferral strategy (SpecCascade [Diff]) often yields better quality trade-offs than the other methods, although the gap to speculative decoding is much smaller.
In Table 2, we report (i) the reduction in latency from different methods when matching the quality of the large model, and (ii) the best quality that each method can deliver without exceeding the latency of the large model (for temperature $T=1$). SpecCascade [OPT] is seen to yield the maximum speed-up and the best quality metrics. The cascading approaches are often seen to fare poorly on both quality and latency metrics, with the exception of WMT, where SeqCascade yields non-trivial speed-ups.
@@ Results.
Table 2: Reduction in latency from different methods when matching the quality of the large model (cols 2–7), and the best quality metric when matching each method yields without exceeding the latency of the large model (cols 8–13). Quality is measured in terms of the BLEU score for WMT and ROUGE-2 for XSum and CNNDM. All results are under temperature sampling with $T=1$.
@@
Figure 2: Plots of quality vs. latency. Each method interleaves T5-small with T5-large (or T5-XL). The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model. See Appendix D for additional plots.
@@ 7 Conclusions
We have proposed new speculative cascading techniques that use a combination of auto-regressive drafting and parallel verification to implement their deferral rule, and shown that they yield better cost-quality trade-offs than standard cascades and speculative decoding. In the future, we wish to replace our plug-in estimators with a router model [18] trained explicitly on ground-truth samples to approximate the optimal rule. We also wish to improve the deferral objective we seek to optimize at each position $t$ (8), and replace it with a global (coupled) deferral objective that takes all prefixes from 1 to $T$ into account. Another useful direction to explore would be to extend our proposal to handle a general cascade with more than two models.
@@ A.1 Proof of Lemma 1
Expanding the loss in (3), we have:
@@ A.2 Proof of Lemma 2
The proof follows straight-forwardly from the results in [44]. Recall from §2 that the lossy speculative decoding procedure of [44] accepts a draft token $x$ with probability:
@@
results in the same acceptance probability (11) and residual distribution (12).
@@ A.3 Proof of Lemma 3
Under a target distribution $\pi_{t}$, the probability of a draft token drawn from $q_{t}$ being is rejected is given by [26]:
@@ A.4 Proof of Lemma 4
Expanding the deferral risk in (8), we have:

### v2

@@ 4.3 Optimal speculative cascade deferral
One can now run the speculative cascading procedure in Algorithm 5 using equation 10 as the deferral rule; the corresponding $\delta(\cdot)$ is listed in Table 1. See §C.2 for a similar derivation for $\ell=\ell_{\log}$.
4.4 Token-specific speculative cascades
The plug-in deferral rules in (5) and (10) decide between the drafter’s distribution $q_{t}(\cdot)$ and the verifier’s distribution $p_{t}(\cdot)$ by comparing their maximum token probabilities. A downside to this approach is that the draft token $x_{t}\sim q_{t}(\cdot)$ may not maximize $q_{t}(\cdot)$. Thus, even when $x_{t}$ is of poor quality, we may end up accepting it because $q_{t}$ happens to be more peaked than $p_{t}$.
To alleviate this problem, we propose the use of token-specific deferral rules $r:\mathscr{V}^{t-1}\times\mathscr{V}\rightarrow\{0,1\}$ that use both the prefix $x_{<t}$ and a candidate token $v$ to provide a binary decision $r(x_{<t},v)\in\{0,1\}$, with 0 indicating that the token is of acceptable quality. We may then construct a target distribution of the following form:
$\pi_{\rm\tt Token}(v)=q_{t}(v)\cdot(1-r(x_{<t},v))+p_{t}(v)\cdot\eta,$ (11)
where $\eta=\sum_{v^{\prime}\in\mathscr{V}}r(x_{<t},v^{\prime})\cdot q_{t}(v^{\prime})$ is a normalizing term chosen to ensure that $\sum_{v}\pi_{\tt Token}(v)=1$. This target distribution closely mimics $q_{t}(\cdot)$ on tokens that the deferral rule $r$ deems to be of acceptable quality, and defers to $p_{t}(\cdot)$ otherwise. One can modify the generic speculative sampling algorithm in Algorithm 4 to use $\pi_{\tt Token}$ as the target distribution, as shown in Algorithm 6 in §D.
To design $r$, we propose a heuristic variant of the Diff rule in equation 4 that compares the expected 0-1 loss from the candidate token $v$ with the expected 0-1 loss from distribution $p_{t}$ (in §D, we discuss deriving a similar variant of the OPT rule in equation 9): color=RoyalBlue!35!white]the reviewer may wonder how these relate to Equations 4 and 9. Hari: added refs to eq 4 and 9.
$r(x_{<t},v)=1\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ 1-\mathbb{P}(v|x_{<t})\,>\,\mathbb{E}_{v\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell_{\text{0-1}}(v,p_{t})\right]+\alpha,$ (12)
for a cost parameter $\alpha$. The following are some simple plug-in approximations to equation 12:
$\hat{r}_{\rm\tt TokenV1}(x_{<t},v)=1$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \textstyle q_{t}(v)\,<\,\max_{v^{\prime}}p_{t}(v^{\prime})-\alpha$ (13)
$\hat{r}_{\rm\tt TokenV2}(x_{<t},v)=1$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \textstyle p_{t}(v)\,<\,\max_{v^{\prime}}p_{t}(v^{\prime})-\alpha$ (14)
$\hat{r}_{\rm\tt TokenV3}(x_{<t},v)=1$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \iff\leavevmode\nobreak\ \leavevmode\nobreak\ \textstyle p_{t}(v)\,<\,\max_{v^{\prime}}p_{t}(v^{\prime})\cdot(1-\alpha),$ (15)
where we approximate $\mathbb{P}(v|x_{<t})$ with either $q_{t}(v)$ or $p_{t}(v)$. Equation 15 is a multiplicative plug-in approximation that has similarities to the rejection criterion used by Leviathan et al. (2023) for lossy speculative greedy decoding, and results in an intuitive target distribution:
$\textstyle\pi_{\tt TokenV3}(v)=q_{t}(v)\cdot\bm{1}\big{(}v\in\textup{Top}_{\alpha}\big{)}\,+\,p_{t}(v)\cdot\sum_{v^{\prime}\notin\textup{Top}_{\alpha}}q_{t}(v^{\prime}),$
where $\textup{Top}_{\alpha}=\{v\in\mathscr{V}:\,p_{t}(v)\,\geq\,\max_{v^{\prime}}p_{t}(v^{\prime})\cdot(1-\alpha)\}$ is the set of top ranked tokens by $p_{t}(\cdot)$. For these top-ranked tokens, $\pi_{\tt TokenV3}$ approximates $q_{t}(\cdot)$; for the rest, it is a re-scaled version of $p_{t}(\cdot)$.
5 Further related work and conclusions
@@
There has been a stream of work on improving the draft generation process in speculative decoding; these include having the drafter and verifier share the same backbone (Stern et al., 2018; Kim et al., 2024; Cai et al., 2024; Monea et al., 2023; Hooper et al., 2023; Zhang et al., 2023; Elhoushi et al., 2024; Liu et al., 2024), using multiple small draft models Chen et al. (2023c); Wang et al. (2024), using tree-structured draft batches (Spector & Re, 2023; Miao et al., 2024), distilling the drafter with the verifier (Zhou et al., 2024), and leveraging multiple sampled draft candidates Sun et al. (2024).color=RoyalBlue!35!white]maybe this can come after Experiments? Hari: the problem with doing this is that we reference this section in the expts when pointing to the target distribution for BiLD!
@@
The work that is most closely related to our specific proposal is the Big Little Decoder (BiLD) (Kim et al., 2023), which can be seen as another lossy variant of speculative decoding (Leviathan et al., 2023; Tran-Thien, 2023; Zhou et al., 2024). BiLD has two phases: a fallback phase, during which the drafter $q$ is run auto-regressively until its maximum predicted probability is sufficiently low; and a rollback phase, during which the verifier $p$ is run in parallel on the prefixes generated by $q$ and rolls back to the point where $D(q,p)>\alpha$, for a metric $D$ that measures discrepancy and threshold $\alpha$. The fallback phase implements Chow’s deferral rule in (2), and allows for the draft window size to vary dynamically based on an estimate of how likely the draft tokens will be accepted; the rollback phase can be seen as a deterministic variant of the rejection sampling algorithm of Leviathan et al. (2023).
@@
An advantage of BiLD over the rejection sampling algorithm in (Leviathan et al., 2023) is the use of Chow’s rule to vary the draft window size. However, the final target distribution it seeks to mimic, $\mathbb{T}_{\textup{BiLD}}(q,p)(v)=\bm{1}(D(q,p)\leq\alpha)\cdot q(v)+\bm{1}(D(q,p)>\alpha)\cdot p(v)$, is an approximation to $p$; specifically, the target distribution $\pi=\mathbb{T}_{\textup{BiLD}}(q,p)$ is chosen to satisfy $D(\pi,p)\leq\alpha$. Hence, in cases where $q$ deviates substantially from $p$, BiLD would choose $p$ as the target distribution, even when $q$ offers better quality on a prefix (where quality can be measured using a suitable loss function). In contrast, our proposed approach in §4 uses speculative decoding to approximate target distributions that seek to optimally cascade between $q$ and $p$. In our experiments, we compare the efficacy of using $\mathbb{T}_{\textup{BiLD}}$ as the target distribution with the target distributions we propose in this paper (see Table 1).
@@
Table 2: Reduction in latency from different methods ($T=1,\gamma=5$) when matching the quality of the large model (cols 2–7), and the best quality metric when matching each method yields without exceeding the latency of the large model (cols 8–13). Quality is measured in terms of the BLEU for WMT and ROUGE-2 for XSum and CNNDM. See Figure 2 for $T=0.5$ and §E.2 for $T=0$.
@@
SpecCascade [Token] 1.85$\times$ 1.18$\times$ 1.89$\times$ 2.50$\times$ 1.40$\times$ 1.89$\times$ 22.50 15.85 12.63 22.70 18.79 12.63
@@
Figure 2: Latency-quality trade-off plots for methods that interleave T5-small with T5-large ($\gamma=5$). Latency is measured relative to that of calling T5-large on all inputs. The horizontal dotted line denotes the quality of T5-large. §E.3E.4 contain more plots with varying temperatures and $\gamma$.
6 Experimental results
We compare our speculative cascading techniques with both sequential cascades and standard speculative decoding on a range of language benchmarks, including translation, reasoning, coding, QA, etc. We evaluate speculative cascades constructed from both the T5 v1.1 family of encoder-decoder models (Raffel et al., 2020a), and Gemma v2 decoder-only models (Team et al., 2024). We construct the cascades with four deferral rules: (i) Chow in (2), (ii) Diff in (5), (iii) OPT in (10), and (iv) the Token-specific rule in (15) (we present results for the V1 and V2 variants in §E.7).
Cascades versus SpecDecode evaluation. Our evaluation protocol is markedly different from the standard evaluation of speculative decoding algorithms, where the goal is to speed up inference with a large model while preserving its output distribution. In contrast, our focus is on trading-off quality for lower inference costs by cascading two models of different sizes. color=RoyalBlue!35!white]the “and” seems like it could be a new sentence. it is not related to the point about deviating from the standard SPEED evaluation protocol? Hari: done! We also do not claim to develop a new state-of-the-art method for fast LM inference. Furthermore, the speculative cascades we design build on the original speculative decoding algorithm Leviathan et al. (2023). While one could potentially also adapt our proposal to other recent variants of speculative decoding (Cai et al., 2024; Li et al., 2024b), these involve a wholly orthogonal suite of techniques to what we propose (such as architectural changes, allowing for multiple drafts, distillation, and so on; see §5).
Baselines. The cascading and speculative decoding methods we compare to include:
  1. (i)
    Sequence-level cascade (Jitkrittum et al., 2023; Gupta et al., 2024) based on sequence-level Chow’s rule in equation 1 (SeqCascade [Chow]).
  2. (ii)
    Token-level cascade outlined in Algorithm 2, with token-level Chow’s rule in equation 2 used for deferral (Chow, 1970; Gupta et al., 2022) (TokenCascade [Chow]).
  3. (iii)
    Lossy speculative decoding described in §2, with both $\beta=1$ (Leviathan et al., 2023; Zhou et al., 2024) (SpecDecode [Lossy]) and $\beta$ tuned using the procedure in Tran-Thien (2023) (Lossy).
  4. (iv)
    Big-Little Decoder approach (Kim et al., 2023), with both the original deterministic version (BiLD), and the variant where we apply Algorithm 4 to the target distribution $\mathbb{T}_{\textup{BiLD}}$ in §5 (BiLD).
[…]
Figure 3: Plots of quality vs. rejection rate for methods that interleave Gemma 2B with Gemma 27B ($\gamma=1$). We use instruction-tuned models; for MBPP we report additional results with pre-trained models. See §E.7 for remaining plots, comparison to (1314) and results on 2B $\rightarrow$ 9B cascades.
In Figure 2, we present plots of quality vs. latency for the different methods. In each case, we vary the lenience parameter $\alpha$, and plot either the BLEU or ROUGE-2 metric as a function of the relative latency to the larger model. For brevity, we include the three main baselines; in §E.5E.6, we compare to SpecDecode [Lossy] (Tran-Thien, 2023) and the original BiLD algorithm Kim et al. (2023). Methods that use speculative execution are considerably faster than sequential token-level cascades (TokenCascade [Chow]), although sequential cascades do have an advantage in the low-latency regimes. This is because unlike speculative approaches, which always call the large model after every $\gamma$ steps, a sequential cascade only invokes the large model when the small model defers. color=RoyalBlue!35!white]the legend is a bit hard to make out. we may consider only including two plots per row. Hari: Did this for Fig 2. For Fig 3, I’ve made the font size bigger. Would you like two figures per row for Fig 3 too? color=RoyalBlue!35!white]further to the above, we may consider the value of including full deferral curves for both T5 and Gemma, versus having a summary like Table 2. Hari: hmm, I think reviewers usually like to see numbers in a tabulated form, especially when it comes to speed-ups
In Table 2, color=RoyalBlue!35!white]caption should clarify this is for T5. Hari: done! we report (i) the reduction in latency from T5 cascades when matching the quality of the large model, and (ii) the best quality that each method can deliver without exceeding the latency of the large model. SpecCascade [Token] often yields the highest speed-up and the best quality metrics, with OPT coming in second. The cascading approaches are often seen to fare poorly on both quality and latency metrics, with the exception of WMT, where SeqCascade yields non-trivial speed-ups.color=cyan]Do we need to briefly remind the reader that BiLD* is different from BiLD implementation? The reason the Token-specific rule fares better than OPT and Diff is because the latter compute their deferral decisions based on which of $q_{t}(\cdot)$ and $p_{t}(\cdot)$ is more peaked; this can be a disadvantage when the sampled token is not close to the distribution mode, which is likely to happen with higher temperatures. As shown in §E.3, with lower temperatures, the gap between these rules diminishes.color=RoyalBlue!35!white]should we also call out the WMT results being sensitive to temperature? Hari: I think this is true of all methods. I’ll be including plots with varying temperatures in the appendix. After I include that, I’ll add a note in the main text.
Few-shot Gemma cascades. color=RoyalBlue!35!white]what about the MBPP results? Hari: TBD To evaluate the Gemma model cascades, we use few-shot prompting with 8 language benchmarks: WMT, CNN/DM, GSM8K, MBPP, SQuAD 2.0, WebQuestions, NaturalQA and TriviaQA; many of these feature in the SpecBench suite (Xia et al., 2024b). Figure 3 presents plots of quality vs. rejection rate with a 2B drafter and 27B verifier for $\gamma=1$. For brevity, we only compare the methods that fare the best in the previous experiments. With the exception of TriviaQA, SpecCascade [Token] is able to both match the 27B’s quality at a lower rejection rate and yield the best overall quality, often better than 27B. Since all three methods use the exact same implementation for speculative execution, a lower rejection rate directly translates to a lower latency.
Interestingly, OPT is not as effective as with T5. We attribute this to the differences in distributions between the two setups: with T5, the maximum token probability served as a good indicator of token accuracy for both $q$ and $p$; with Gemma, we expect the large model to have a closer alignment with the ground-truth distribution (due to it being several billion parameters apart from the smaller model), and hence using the large model probabilities to measure confidence for both the small and large model (15) yields better trade-offs than comparing the modes from the two model distributions.
@@ 7 Conclusions
We have proposed new speculative cascading techniques that use a combination of auto-regressive drafting and parallel verification to implement their deferral rule, and shown that they yield better cost-quality trade-offs than standard cascades and speculative decoding. A limitation of our approach is that while it offers a higher throughput, it also incurs a higher total compute cost compared to sequential cascades. In the future, we wish to replace our plug-in estimators with a router model (Gupta et al., 2024) trained on ground-truth samples, to improve the local deferral objective at each position $t$ (8) with a global objective, and to extend our proposal to more than two models.
@@ A.1 Proof of Lemma 1
Expanding the loss in equation 3, we have:
@@ A.2 Proof of Lemma 2
The proof follows straight-forwardly from the results in (Tran-Thien, 2023). Recall from §2 that the lossy speculative decoding procedure of (Tran-Thien, 2023) accepts a draft token $x$ with probability:
@@
results in the same acceptance probability equation 16 and residual distribution equation 17.
@@ A.3 Proof of Lemma 3
Under a target distribution $\pi_{t}$, the probability of a draft token drawn from $q_{t}$ being is rejected is given by (Leviathan et al., 2023):
@@ A.4 Proof of Lemma 4
Expanding the deferral risk in equation 8, we have: